# Haptic bilateral teleoperation system for free-hand dental procedures

### v2

@@ III-B Robot interaction control
In this section, we first formalize the contact model between the robot and the environment (Section III-B1), and subsequently, we present the free-hand strategy for limiting the interaction forces (Section III-B2).
@@ III-B1 Admittance model
Let ${\bf x}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)=\begin{bmatrix}{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)^{\top}&{\bf q}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)^{\top}\end{bmatrix}^{\top}$ be the desired pose and $\mathcal{R}_{c}$ the compliant frame specified by the quantities ${\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}(k)$, $\dot{\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}(k)$, $\ddot{\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}(k)$ and ${\bf q}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}(k)$. We do not consider compliance along the rotational axes, therefore, we assume ${\bf q}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}(k)={\bf q}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$. When the FR interacts with the environment, $\mathcal{R}_{c}$ can be completely defined from ${\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$, $\dot{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$ and $\ddot{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$, and the contact forces ${\bf f}^{\mathcal{R}_{e}}(k)\in\mathbb{R}^{3\times 1}$, by integrating the mechanical impedance equation:
${\bf M}_{d}\ddot{\tilde{{\bf p}}}^{\mathcal{R}_{e}}(k)+{\bf K}_{D}\dot{\tilde{{\bf p}}}^{\mathcal{R}_{e}}(k)+{\bf K}_{P}\tilde{{\bf p}}^{\mathcal{R}_{e}}(k)=-\tilde{\bf f}^{\mathcal{R}_{e}}(k),$ (24)
where ${\bf M}_{d},{\bf K}_{D},{\bf K}_{P}$ are the mass, damping, and stiffness diagonal matrices, respectively, of proper size, which are used to impose specific second-order mass-spring-damper dynamics; $\tilde{\bf p}^{\mathcal{R}_{e}}(k)={\bf p}^{\mathcal{R}_{e}}_{\mathcal{R}_{ed}}(k)-{\bf p}^{\mathcal{R}_{e}}_{\mathcal{R}_{c}}(k)$ is the position error and $\dot{\tilde{\mathbf{p}}}^{\mathcal{R}_{e}}(k)$ and $\ddot{\tilde{\mathbf{p}}}^{\mathcal{R}_{e}}(k)$ its time derivatives; $\tilde{{\bf f}}^{\mathcal{R}_{e}}(k)={\bf f}_{d}^{\mathcal{R}_{e}}(k)-{\bf f}_{e}^{\mathcal{R}_{e}}(k)$ is the force tracking error, defined as the difference between the desired force ${\bf f}_{d}^{\mathcal{R}_{e}}(k)$ and the exerted force ${\bf f}_{e}^{\mathcal{R}_{e}}(k)$, which is equal and opposite to the measured force. We take ${\bf f}_{d}^{\mathcal{R}_{e}}(k)={\bf 0},\forall k\in\mathbb{N}_{0}$ to make the robot behave compliantly with the environment and $\dot{\bf p}^{\mathcal{R}_{e}}_{\mathcal{R}_{ed}}(k)=\ddot{\bf p}^{\mathcal{R}_{e}}_{\mathcal{R}_{ed}}(k)={\bf 0},\forall k\in\mathbb{N}_{0}$ to reach a complete stop of the FR’s end-effector upon reaching the target position, effectively preventing any collisions caused by further movement. From (24), the compliant position reference to be commanded to the robot is obtained as:
${\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{b}}(k)={\bf p}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)+{\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k){\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}(k).$ (25)
@@ III-B2 Free-hand force limitation strategy
Assuming that the environment can be approximated by a linear spring model, the exerted force at the steady state [47] is:
${\bf f}_{ss}^{\mathcal{R}_{e}}={\bf K}_{eq}\left({\bf p}_{\mathcal{E}_{r}}^{\mathcal{R}_{e}}-{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\right),$ (26)
where ${\bf K}_{eq}=({\bf K}_{P}{\bf K}_{e})({\bf K}_{P}+{\bf K}_{e})^{-1}$ is the equivalent stiffness of the admittance stiffness ${\bf K}_{P}$ and environmental stiffness ${\bf K}_{e}$ [48], and ${\bf p}_{\mathcal{E}_{r}}^{\mathcal{R}_{e}}$ is the environment rest position. From (26), it follows that interaction forces can be limited through a suitable reference ${\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}$ provided to the controller. Accurate limitation requires perfect knowledge of the environment that is particularly challenging to achieve, especially in a medical environment, in which the FR interacts with objects of varying stiffness and non-uniform geometry. Therefore, in the proposed strategy, the position reference:
${\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)=\left[{\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)\right]^{\top}{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)-\left[{\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)\right]^{\top}{\bf p}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k),$ (27)
@@
$\underline{{\bf p}}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)=\frac{1}{1+K_{s}\left\lVert{\bf f}_{e}^{\mathcal{R}_{e}}(k)\right\rVert}{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k),$ (28)
@@
where $K_{s}\in\mathbb{R}^{+}$ is a proportional gain representing a trade-off between stability and responsiveness of the scaling action. Indeed, $K_{s}$ close to 00 reduces sudden variations of $\underline{{\bf p}}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$, even in presence of noisy force measures, but provides less force-limiting action. On the other hand, $K_{s}$ significantly greater than 00 makes the scaling action more responsive, but may yield oscillations in the presence of noisy measurements. Therefore, to increase the system’s robustness to noisy forces, exhibiting high-frequency variations, ${\bf f}_{e}^{\mathcal{R}_{e}}$ is filtered online as done for positions in (6). We note that the strategy in (28) does not allow forces to be bounded below a predefined limit. However, it prevents contact forces from growing indefinitely, ensuring a limitation even in the transients of the interaction and for any orientation of the end-effector, without relying on the knowledge of the dental environment’s complex geometry.
@@ III-C Haptic feedback controller
In performing medical procedures, the operator relies heavily on the sense of touch. Therefore, in a HBTS providing realistic haptic sensations, it becomes critical to the success of the procedures. In impedance systems, the interaction is captured by both the contact force ${\bf f}_{e}^{\mathcal{R}_{e}}$ and the position error $\tilde{\bf p}^{\mathcal{R}_{e}}$, the latter being characterized by slower transients than the former, which exhibits a step dynamics. As proposed in [49], replacing ${\bf f}_{e}^{\mathcal{R}_{e}}$ with $\tilde{\bf p}^{\mathcal{R}_{e}}$ to generate the force feedback makes each contact state transparent to the operator, preventing them from having to deal with sudden force changes they are not used to from previous training and fieldworks. Therefore, the proposed HFC defines a second (virtual) impedance model, generating the force feedback from the motion error of the admittance controller:
@@
where ${\bf K}_{h}$ and ${\bf D}_{h}$ respectively are virtual stiffness and damping diagonal matrices of the proper size. The position error $\tilde{{\bf p}}^{\mathcal{H}_{b}}(k)$ and its time derivative $\dot{\tilde{{\bf p}}}^{\mathcal{H}_{b}}(k)$ are defined, respectively, as:
$\tilde{{\bf p}}^{\mathcal{H}_{b}}(k)$ $={\bf R}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\left[{\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)\right]^{\top}\tilde{{\bf p}}^{\mathcal{R}_{e}}(k),$ (30)
$\dot{\tilde{{\bf p}}}^{\mathcal{H}_{b}}(k)$ $={\bf R}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\left[{\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)\right]^{\top}\dot{\tilde{{\bf p}}}^{\mathcal{R}_{e}}(k).$ (31)
@@
We note that, as in (17), the rotation matrix ${\bf R}_{\bf z}^{\mathcal{H}_{e}}(\varphi(k))$ is used to provide the force feedback correspondence with the interaction observed in the camera view. With regard to the ${\bf K}_{h}$ and ${\bf D}_{h}$ parameters of the virtual system in (29), they are tuned to provide a natural force feedback to the human operator while the FR interacts with the environment, decoupling it from the admittance control dynamics. Specifically, the stiffness matrix ${\bf K}_{h}$ can be tuned to achieve the desired transient response, while the damping matrix ${\bf D}_{h}$ can be tuned to mitigate vibrations. We envisage the possibility that the generated force feedback may exceed the range of forces that the HD device can exert. Therefore, the HFC saturates the norm of the force feedback $\lVert{\bf f}_{h}^{\mathcal{H}_{b}}(k)\rVert$ to the maximum exertable force $f_{MAX}^{\mathcal{H}_{b}}\in\mathbb{R}^{+}$:
@@
$\underline{\mathbf{f}}_{h}^{\mathcal{H}_{b}}(k)=\begin{cases}\mathbf{f}_{h}^{\mathcal{H}_{b}}(k),&\text{if}~{}\lVert\mathbf{f}_{h}^{\mathcal{H}_{b}}(k)\rVert<f_{MAX}^{\mathcal{H}_{b}};\\ f_{MAX}^{\mathcal{H}_{b}}\dfrac{\mathbf{f}_{h}^{\mathcal{H}_{b}}(k)}{\lVert\mathbf{f}_{h}^{\mathcal{H}_{b}}(k)\rVert},&\text{otherwise}.\end{cases}$ (32)
@@
Finally, given the slow dynamics of $\tilde{{\bf p}}^{\mathcal{R}_{e}}$, the human operator may perceive residual force feedback even when the robot is no longer interacting. To address this, we introduce a force dead-band $f_{db}\in\mathbb{R}^{+}$ to prevent the HD from reacting to such residuals:
@@
$\underline{{\bf f}}_{h}^{\mathcal{H}_{b}}(k)=\begin{cases}\underline{{\bf f}}_{h}^{\mathcal{H}_{b}}(k),&\text{if}~{}\left\lVert{\bf f}_{e}^{\mathcal{R}_{e}}(k)\right\rVert>f_{db};\\ {\bf 0},&\text{otherwise}.\end{cases}$ (33)
IV Experimental validation

### v3

@@ III-B Robot interaction control
In this section, we first formalize the contact model between the robot and the environment (Section III-B1), and subsequently, we present the free-hand strategy for limiting the interaction forces (Section III-B2) and the other safety mechanisms embedded in the controller (Section III-B3).
@@ III-B1 Admittance model
Let ${\bf x}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$ be the desired pose and $\mathcal{R}_{c}$ the compliant frame specified by the quantities ${\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{ed}}(k)$, ${\bf q}_{\mathcal{R}_{c}}^{\mathcal{R}_{ed}}(k)$, ${\bf v}_{\mathcal{R}_{c}}^{\mathcal{R}_{ed}}(k)$, $\dot{\bf v}_{\mathcal{R}_{c}}^{\mathcal{R}_{ed}}(k)$. When the FR interacts with the environment, $\mathcal{R}_{c}$ can be completely defined in $\mathcal{R}_{ed}$ by integrating the mechanical impedance equation:
${\bf M}_{d}\dot{{\bf v}}^{\mathcal{R}_{ed}}_{\mathcal{R}_{c}}(k)+{\bf K}_{D}{{\bf v}}^{\mathcal{R}_{ed}}_{\mathcal{R}_{c}}(k)+{\bf K}_{P}{\bf x}^{\mathcal{R}_{ed}}_{\mathcal{R}_{c}}(k)=-\tilde{\bf h}^{\mathcal{R}_{ed}}(k),$ (17)
where ${\bf M}_{d},{\bf K}_{D},{\bf K}_{P}\in\mathbb{R}^{6\times 6}$ are the mass, damping, and stiffness diagonal matrices, respectively, which are used to impose specific second-order mass-spring-damper dynamics; $\tilde{{\bf h}}^{\mathcal{R}_{ed}}(k)={\bf h}_{d}^{\mathcal{R}_{ed}}(k)-{\bf h}_{e}^{\mathcal{R}_{ed}}(k)$ is the wrench tracking error, defined as the difference between the desired wrench ${\bf h}_{d}^{\mathcal{R}_{ed}}(k)$ and the exerted one ${\bf h}_{e}^{\mathcal{R}_{ed}}(k)$, which is equal and opposite to the measured wrench. We take ${\bf h}_{d}^{\mathcal{R}_{ed}}(k)={\bf 0},\forall k\in\mathbb{N}_{0}$ to make the robot behave compliantly with the environment.
@@ III-B2 Free-hand force limitation strategy
In an admittance-controlled system, the interaction wrenches emerge from the combination of the admittance parameters, the mechanical properties of the environment, and the reference provided to the controller [20]. It follows that interaction forces can be limited through a suitable reference ${\bf x}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}$. Accurate limitation requires perfect knowledge of the environment that is particularly challenging to achieve, especially in a medical environment, in which the FR interacts with surfaces of varying stiffness and non-uniform geometry. Therefore, in the proposed strategy, the position reference
${\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)={\bf R}_{\mathcal{R}_{b}}^{\mathcal{R}_{e}}(k){\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)-{\bf R}_{\mathcal{R}_{b}}^{\mathcal{R}_{e}}(k){\bf p}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k),$ (18)
@@
$\underline{{\bf p}}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)=\frac{1}{1+K_{s}\lVert{\bf f}_{e}^{\mathcal{R}_{e}}(k)\rVert}{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k),$ (19)
@@
where $K_{s}\in\mathbb{R}_{+}$ is a proportional gain representing a trade-off between smoothness and responsiveness of the scaling action. Indeed, $K_{s}$ close to 0 reduces sudden variations of $\underline{{\bf p}}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$, even in presence of noisy force measures, but provides less force-limiting action. On the other hand, $K_{s}$ significantly greater than 0 makes the scaling action more responsive, but may yield oscillations in the presence of noisy measurements. Therefore, to increase the system’s robustness to noisy forces, exhibiting high-frequency variations, ${\bf f}_{e}^{\mathcal{R}_{e}}$ is filtered online as done for positions in (4). We note that the strategy in (19) does not allow forces to be bounded below a predefined limit. However, it prevents contact forces from growing indefinitely, ensuring a limitation even in the transients of the interaction and for any orientation of the end-effector, without relying on the knowledge of the dental environment’s complex geometry.
From a control theory perspective, it is appropriate to draw an analogy between the proposed strategy and passivity-based methods, such as TDPA [7], as both rely on modulating the reference provided to the downstream controller through an adaptive damping mechanism. However, a substantial difference exists in terms of control objectives. While passivity-based approaches provide an event-triggered modulation of the reference in response to violations of the energy balance, with the explicit goal of guaranteeing system passivity, the proposed strategy implements a continuous, non-energizing, force-driven reference modulation aimed at reducing interaction forces, without relying on explicit energy estimation. As a consequence, its attenuative and smoothly varying action on the reference, resulting from an appropriate choice of the gain $K_{s}$ and of the force filtering window, does not inject energy into the system, nor does it alter the underlying dynamics, in contrast to state-of-the-art variable admittance–based approaches [20, 10], which explicitly modify the controller dynamics and typically rely on explicit knowledge of the environment [48]. As a result, the stability properties of the admittance-controlled interaction remain unaffected. Moreover, as discussed above, passivity-based control addresses an orthogonal objective to the proposed approach, and can therefore be combined with the eye–hand coordination and force limitation strategies to guarantee stability even in the presence of considerable communication delays. We consider this scenario in Section IV-B.
III-B3 Safety mechanisms and emergency stop protocols
From (17) and (19), the compliant pose ${\bf x}_{\mathcal{R}_{c}}^{\mathcal{R}_{b}}(k)$ to be commanded to the robot is obtained. Specifically,
${\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{b}}(k)$ $=\underline{{\bf p}}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)+{\bf R}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k){\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{ed}}(k),$ (20)
${\bf q}_{\mathcal{R}_{c}}^{\mathcal{R}_{b}}(k)$ $={\bf q}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k){\bf q}_{\mathcal{R}_{c}}^{\mathcal{R}_{ed}}(k),$ (21)
where ${\bf x}^{\mathcal{R}_{ed}}_{\mathcal{R}_{c}}$, taking the semantics of a pose error, is obtained by time-integrating $\dot{\bf v}^{\mathcal{R}_{ed}}_{\mathcal{R}_{c}}$ from (17) twice. In addition to its compliant interaction properties and the embedded force limitation strategy, the controller incorporates two safety mechanisms and implements an emergency stop protocol. Regarding the safety mechanisms, the first ensures the stability of the FR in the absence of updated references from the EHCC ${\bf x}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$, such as in the event of a communication loss, by continuing to track the last valid one ${\bf x}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k-1)$. The second mechanism prevents abrupt motions of the FR by commanding ${\bf x}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)$ whenever the generated command ${\bf x}_{\mathcal{R}_{c}}^{\mathcal{R}_{b}}(k)$ deviates excessively, either in translation or orientation, based on translational and rotational safety thresholds that can be configured by the operator according to the task requirements. With regard to the emergency stop, it focuses on the monitoring of the interaction forces. Whenever $\lVert{\bf f}^{\mathcal{R}_{e}}\rVert(k)$ exceeds a predefined safety threshold $f_{th}$, the controller forwards ${\bf x}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)$ as the new reference while continuing to compute the admittance control law. In this way, abrupt stopping is avoided, while still allowing the interaction forces to decrease progressively.
@@ III-C Haptic feedback controller
In performing medical procedures, the operator relies heavily on the sense of touch. Therefore, in a HBTS providing realistic haptic sensations, it becomes critical to the success of the procedures. In impedance systems, the interaction is captured by both the contact force ${\bf f}_{e}^{\mathcal{R}_{e}}$ and the position error $\tilde{\bf p}^{\mathcal{R}_{e}}$, the latter being characterized by slower transients than the former, which exhibits a step dynamics. As proposed in [41], replacing ${\bf f}_{e}^{\mathcal{R}_{e}}$ with $\tilde{\bf p}^{\mathcal{R}_{e}}={\bf p}^{\mathcal{R}_{e}}_{\mathcal{R}_{ed}}-{\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}$ to generate the force feedback makes each contact state transparent to the operator, preventing them from having to deal with sudden force changes they are not used to from previous training and fieldworks. Therefore, the proposed HFC defines a second (virtual) impedance model, generating the force feedback from the motion error of the admittance controller:
@@
where ${\bf K}_{h}\in\mathbb{R}^{3\times 3}$ and ${\bf D}_{h}\in\mathbb{R}^{3\times 3}$ respectively are virtual stiffness and damping diagonal matrices, where
$\begin{bmatrix}\tilde{{\bf p}}^{\mathcal{H}_{b}}(k)\\ \dot{\tilde{{\bf p}}}^{\mathcal{H}_{b}}(k)\end{bmatrix}=\Bigl({\bf I}_{2}\otimes\bigl({\bf R}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k){\bf R}_{\mathcal{R}_{e}}^{\mathcal{H}_{e}}(k)\bigr)\Bigr)\begin{bmatrix}\tilde{{\bf p}}^{\mathcal{R}_{e}}(k)\\ \dot{\tilde{{\bf p}}}^{\mathcal{R}_{e}}(k)\end{bmatrix}$. (23)
@@
We note that, as in (9), the rotation matrix ${\bf R}_{\bf z}^{\mathcal{R}_{e}}(k)$ (or equivalently ${\bf R}_{\bf z}^{\mathcal{H}_{e}}(k)$) in ${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)$ is used to provide the force feedback correspondence with the interaction observed in the camera view. This approach is particularly suited to the considered application, as it guarantees stability while providing the operator with a natural and intuitive tactile experience that closely resembles manual execution, especially in light of Remark 2. Within this framework, the parameters ${\bf K}_{h}$ and ${\bf D}_{h}$ of the virtual system in (22) are tuned to shape the perceived interaction, with the stiffness matrix ${\bf K}_{h}$ governing the transient response and the damping matrix ${\bf D}_{h}$ contributing to vibration attenuation. Moreover, the use of a second virtual impedance model to generate the force feedback moderates the influence of transient force spikes associated with sudden contacts, while preserving informative and natural haptic feedback. We envisage the possibility that the generated feedback may exceed the range of forces that the HD device can exert. Therefore, the HFC saturates the norm of the force feedback $\lVert{\bf f}_{h}^{\mathcal{H}_{b}}(k)\rVert$ to the maximum exertable force $f_{\text{max}}^{\mathcal{H}_{b}}\in\mathbb{R}_{+}$:
@@
$\underline{\mathbf{f}}_{h}^{\mathcal{H}_{b}}(k)=\begin{cases}\mathbf{f}_{h}^{\mathcal{H}_{b}}(k),&\text{if}~\lVert\mathbf{f}_{h}^{\mathcal{H}_{b}}(k)\rVert<f_{\text{max}}^{\mathcal{H}_{b}};\\ f_{\text{max}}^{\mathcal{H}_{b}}\dfrac{\mathbf{f}_{h}^{\mathcal{H}_{b}}(k)}{\lVert\mathbf{f}_{h}^{\mathcal{H}_{b}}(k)\rVert},&\text{otherwise}.\end{cases}$ (24)
@@
Finally, given the slow dynamics of $\tilde{{\bf p}}^{\mathcal{R}_{e}}$, the operator may perceive residual force even when the robot is no longer interacting. To address this, we introduce a force dead-band $f_{db}\in\mathbb{R}_{+}$ to prevent the HD from reacting to such residuals:
@@
$\overline{{\bf f}}_{h}^{\mathcal{H}_{b}}(k)=\begin{cases}\underline{{\bf f}}_{h}^{\mathcal{H}_{b}}(k),&\text{if}~\bigl\lVert{\bf f}_{e}^{\mathcal{R}_{e}}(k)\bigr\rVert>f_{db};\\ {\bf 0},&\text{otherwise}.\end{cases}$ (25)
III-D Integrated mechanical end-effector
In order to successfully perform any free-hand dental procedure, the operator relies on three basic factors: (a) the procedure-specific clinical tool; (b) the surgical field view, which may require external illumination and tools to show obstructed or poorly lit areas, e.g., a mirror; (c) the sense of touch to ensure the right pressure to be exerted on the area of interest, in order to achieve a safe and effective procedure. In this work, we design an integrated mechanical end-effector that can provide all the tools and sensory capabilities required to perform free-hand dental procedures.
Figure 6: Integrated mechanical end-effector for free-hand procedures. Visible in the left side of the figure is also one of the dental models used in the experiments.
With reference to Figure 6, the designed end-effector consists of five main components: (1) a 3D-printed mechanical interface connecting the robot’s flange with the force/torque sensor; (2) a force/torque sensor measuring contact forces; (3) a 3D-printed tool socket supporting the clinical tool and the endoscopic camera: the interface consists of a rigid structure to enable proper transmission of forces from the tool to the sensor, a circular hole, adjustable via a bolt-and-nut mechanism, to accommodate tools of different sizes, and a movable camera-support ring, allowing for adjustments of the view and light beam angle; (4) an high definition endoscopic camera, equipped with LED illumination to ensure good visibility even in poorly lit areas; (5) a clinical tool for performing the procedure of interest: the tool is placed in the supporting socket after being wrapped with friction tape to ensure good mechanical compliance, emulating the skin friction between the operator’s hand and the tool; the tool used for the chosen use case is a scaler.
Although the current implementation does not fulfill yet all clinical requirements, such a design is compatible with future clinical adoption, as the end-effector is inherently compatible with standard sterilization and infection-control procedures commonly adopted in dentistry. Indeed, clinical instruments are intrinsically designed to withstand autoclave sterilization, and the literature reports the availability of sterilizable endoscopic cameras [9], as well as autoclavable or low-temperature–sterilizable materials, including medical-grade 3D-printed polymers, which would enable full sterilization of the entire assembly [50]. Moreover, the tool socket has been designed with smooth external surfaces and a reduced number of crevices, thereby facilitating cleaning and reprocessing. In addition, the socket enables rapid tool replacement by simple extraction and insertion of the clinical instrument, without requiring disassembly of the end-effector or disconnection of sensing and vision components. Components upstream of the end-effector, such as the force/torque sensor and the robot flange, which are not compatible with autoclave sterilization, can be isolated using sterile drapes or disposable sterile barriers, as commonly adopted in medical robotic systems [60]. With respect to aerosol management, a critical concern in dental procedures such as SRP, the proposed design does not interfere with conventional mitigation strategies. In a clinical deployment, aerosol generation could be managed through standard high-volume evacuation systems and saliva ejectors, without requiring any modification to the design.
IV Technical validation